# Ruiyan Han2∗  Zhen Fang1  Xinyu Sun2∗  Yuchen Ma2  Ziheng Wang2  Yu Zeng122footnotemark: 2 Zehui Chen1   Lin Chen1   Wenxuan Huang3,4   Weijie Xu5   Yi Cao6   Feng Zhao1 1

### v1

@@ 1 Introduction
The realization of Artificial General Intelligence (AGI) requires a tight synergy between comprehension and generation, wherein comprehension enables the internalization of knowledge and generation allows its coherent and expressive externalization. By integrating multiple modalities into a shared representational space, Unified Multimodal Models (UMMs) bagel; blip3o; xie2025show naturally couple comprehension and generation as two complementary phases of a unified cognitive process, supporting both knowledge grounding and coherent reasoning.
@@
Across extensive experiments, we find that our model achieves reliable self-improvement without heuristic reward engineering, curriculum design, or external supervision. Compared with prior self-improvement approaches srum and methods that depend on external guidance, our approach learns from internally generated training signals, generalizes well, and remains stable under out-of-distribution (OOD) conditions. These results support the effectiveness of a fully self-contained learning paradigm.
@@ 2 Related Work
UMMs aim to unify cross-modal understanding and generation, yet strong understanding often fails to yield equally strong native generation. Existing designs fall into two paradigms: pure autoregressive models that jointly predict text and visual tokens over interleaved sequences  chen2025janus; cui2025emu3; tong2025metamorph) and hybrid models that combine autoregressive language modeling with diffusion-based image synthesis, either within a unified backbone  xie2024show; zhao2024monoformer) or via modular routing and sparse experts shi2024lmfusion; liang2024mixture; deng2025emerging), with related guidance schemes such as Diffusion Forcing chen2024diffusion. Beyond architecture, self-improvement methods convert self-generated signals into training objectives yu2025guided; zhou2024calibrated; wang2025unified; for UMMs, SRUM derives internal rewards from understanding srum, and UniRL jointly optimizes understanding and generation unirl. However, most pipelines depend on auxiliary components or task-specific feedback, limiting scalability and generalization.
@@
Multi-agent systems decompose reasoning through role specialization and interaction, enabling solution diversity and cross-verification, but often incur high coordination cost and brittle verification chen2024agentverse; liang2024encouraging; cemri2025multi. In parallel, LLM self-improvement converts self-generated tasks and evaluations into training signals, supporting zero-data learning via self-play and self-rewarding mechanisms silver2017mastering; huang2025r; zhao2025absolute; yuan2024self. Unified Multimodal Models (UMMs) naturally unify understanding and generation within a single parameter space, making them particularly well-suited for lightweight role instantiation and fully model-driven self-improvement without external supervision.
@@ 3.1 Motivation
Just as a child who associates the word “apple” with the fruit can spontaneously name it upon seeing it, cognitive symmetry blanco2018unconscious enables a bidirectional mapping between internal concepts and external expressions. This alignment is reminiscent of escaping Plato’s Cave: true intelligence must move beyond observing surface data to mastering the reciprocal relationship between an appearance and its underlying source.
@@
Figure 3: Results of BAGEL bagel and GPT-4o gpt4o on four understanding benchmarks. For Omini-RewardBench jin2025omni and MMRB2 hu2025multimodal, we evaluate the T2I task. Performances are normalized with GPT-4 achiam2023gpt results for better visualization.
@@ 3.3.1 Stage 1: Self Multi-Agent Sampling
LLMs are naturally suited for self-play in multitask settings radford2019language. For UMMs, interleaved multimodal inputs and functional diversity allow prompting, generation, and judgement to coexist within a shared model, enabling role-conditioned behaviors under different prompts. We leverage this property to functionalize a single UMM into collaborative roles, bridging the comprehension–generation gap through internal synergy.
@@
The proposer is designed to generate a diverse set of challenging prompts for the unified multimodal model, which are subsequently used to produce training images. To this end, inspired by LAION-5B laion-5b and COYO-700M coyo-700m, we partition all T2I task prompts into ten categories and designed fine-grained generation rules for each category. Next, we prompt UMM to generate an initial batch of prompts and act as the judge to select the best candidate for subsequent iterations. Leveraging the strong in-context learning (ICL) capabilities of LLMs dong2024survey, the initial example serves as a few-shot demonstration to guide the generation of subsequent prompts. To further enhance diversity, we introduce a dynamic seeding mechanism. After generating a predefined number of prompts, several examples are sampled from the prompt library for evaluation and then used to construct new demonstrations that guide the next round of prompt generation. Compared with prior approaches that either directly rely on training set srum or employ external models for prompt construction unirl, our method requires no external data and generates more diverse prompts, thereby improving generalization.
@@
The solver is responsible for producing a diverse set of outputs in response to the prompts generated by the proposer. Therefore, we encourage the UMM to generate images under random seeds and different hyperparameters. Following DeepSeek-R1 deepseek-r1, we perform 8 rollouts per prompt to strike a favorable trade-off between sample quality, diversity, and computational efficiency.
@@
Previous work has relied on heuristic reward functions based on keywords unirl or on powerful external models to provide dense reward maps srum. Such reward judges depend heavily on parameter tuning and the performance of external models, which varies across tasks, thereby severely limiting the generalization of self-improvement. As illustrated in Fig. 3, UMMs exhibit strong reward modeling capabilities. Thus, we formulate reward evaluation for all T2I tasks using discrete scores ranging from 0 to 10, following a widely adopted LLM-as-a-judge paradigm radford2019language; kim2023prometheus. To further enhance judgement quality, we transfer generation reward models deepseek-grm, which have demonstrated strong potential in LLMs, to T2I evaluation. Specifically, we design task-specific rubrics for each category and encourage the model to explicitly articulate its reasoning before producing the final score.
@@ 3.3.2 Stage2: Cognitive Pattern Reconstruction
Through self multi-agent rejection sampling using the Proposer–Solver–Judge pipeline, we obtain a batch of high-quality prompt–image pairs. While these pairs reflect a mapping from abstract conceptual spaces to high-dimensional visual manifolds, directly optimizing this cross-domain alignment remains stochastic and inefficient, often leading to mode collapse chen2025t2i; wang2024div. To move beyond this "black-box" optimization, we draw inspiration from metacognitive theory dunlosky2008metacognition, which identifies monitoring, evaluation, and regulation as the pillars of robust learning. Based on this insight, we propose a tripartite data architecture that reclaims and structures the overlooked trajectories from the self-play cycle. By replaying these latent interactions as explicit caption, judgement, and reflection patterns, we respectively ground abstract concepts in visual features, provide evaluative signals, and encode self-correction processes. This design transforms the previously discarded internal "inner monologue" into a structured supervisory signal, fostering cognitive symmetry without external intervention.
@@
Reflection Inspired by Reflexion shinn2023reflexion, this pattern introduces iterative regulation to enhance the model’s capacity for self-evolution. Leveraging the Solver’s multiple rollouts $\{I_{1},\dots,I_{n}\}$, we utilize the rewards assigned by the Judge to identify pairs of contrasting quality, specifically selecting a high-reward "winning" image $I^{*}$ and a lower-reward "losing" image $I_{lose}$ from the same prompt. We then construct reflection trajectories formulated as $\pi_{\theta}(I^{*}\mid T,I_{lose},J)$, which explicitly encode the transition from suboptimal states to superior ones. By learning to transform the lower-quality manifestation $I_{lose}$ into its optimized counterpart $I^{*}$, the model internalizes a mechanism for self-correcting generative errors, effectively mitigating mode collapse without the need for external supervision.
@@ 3.4 UniCycle
Based on TIIF wei2025tiifbenchdoest2imodel, we generate QA pairs to probe instruction-implied attributes grounded in the generated image, extending the original TIIF benchmark from the T2I setting to the Text-to-Image-to-Text (T2I2T) setting. After annotation, we obtain 1,401 TIIF-style instances that cover more than ten task categories and span multiple question formats, including multiple-choice, binary (yes/no), and open-ended questions.
@@ 4.1 Experiment Setup
We adopt BAGEL bagel as the base model for our main experiments. The Proposer generates 5,000 prompts, then the Solver rolls out 8 times for each prompt. Training is conducted for 600 steps on 8 NVIDIA H800 GPUs for about 7 hours with a constant learning rate of $1\times 10^{-5}$. Additional details are provided in the Appendix  A.
@@
To validate our method, we compare it against three categories of approaches. First, we consider baseline models, including T2I frameworks: SD3 Medium SD3, FLUX.1-dev flux and unified multimodal models: Janus-Pro januspro2025, Show-o2 xie2025show, BLIP3-o blip3o, UniGen tian2025unigen, TwiG guo2025thinking and T2I-R1 jiang2025t2i.
@@
Regarding evaluation, we focus on TIIF wei2025tiifbenchdoest2imodel, WISE niu2025wise, OneIG-EN chang2025oneigbenchomnidimensionalnuancedevaluation, CompBench kil2024compbench, DPG hu2024ella, and Geneval ghosh2023geneval to assess generation performance. To evaluate understanding, we further report results on standard benchmarks including MME fu2023mme, MMB liu2024mmbench, MMMU-val yue2024mmmu, MMVP tong2024eyes, and MMStar chen2024we.
@@ 4.3.2 Model Architecture
We first demonstrate the effectiveness of our method on BAGEL, where understanding and generation components are decoupled. To evaluate its generalization to tightly coupled architectures, we conduct a base model ablation on the purely autoregressive Janus-Pro-7b januspro2025. Tab. 4 shows that our method improves Janus by +3.2 on TIIF, +7.0 on WISE, and +4.7 on OneIG-EN, with the most pronounced gain on the knowledge-intensive WISE benchmark. This suggests that the proposed approach enhances knowledge expression by guiding generation through improved understanding, a mechanism that generalizes across different model architectures.
@@ 4.3.3 Scaling Law for UniCorn
Scaling laws guide architectural design and optimization kaplan2020scaling; chen2024agent, but prior methods scale poorly due to reliance on external models or fixed prompts. In contrast, UniCorn achieves scalable self-improvement through unbounded self-sampling and efficient Cognitive Pattern Reconstruction. To explicitly assess scalability, we conduct scale-up experiments by varying the amount of self-generated data across {1k, 5k, 8k, 10k, 20k}.
@@
As shown in Fig. 7, with only 1K training samples, our method already surpasses RecA on TIIF. As the data scale increases, the model’s generative performance continues to improve, with more pronounced gains on long-prompt generation; notably, with just 5k samples, it outperforms IRG interleaving-reasoning trained on 30k GPT-4o distilled data as well as the strong closed-source model DALL·E 3 betker2023improving. These results reveal a favorable scaling regime in which self-generated data alone suffices to drive continual and efficient improvements in generative capability.
@@ 4.4 Analysis
For self-play assessment, we use Qwen3-VL-235B-A22B-Instruct yang2025qwen3 for data construction (UniCorn*). As shown in Tab. 4, employing stronger proposers/judges yields diminishing returns, where high costs and training time outweigh performance gains. This likely stems from the difficulty of fitting high-entropy teacher distributions, increasing latency without proportional information gain. We then compare UniCorn with four unified models on UniCycle (Judge: Qwen3-235B-A22B yang2025qwen3). UniCycle requires both generation and understanding, reducing task bias and evaluating the model’s self-reflection, thus signaling comprehensive multimodal intelligence.
@@ 5 Conclusion
Content
Due to space limitations, we present additional details, along with quantitative and qualitative results of our UniCorn, in the appendix. The outline is provided below.
  • A. Additional Details (Appendix A)
    • Details of Training Data
    • Training Setup
    • Details of T2I Benchmarks
  • B. More Related Work (Appendix B)
    • Unified Multimodal Models
    • Self-Improvement for LLMs
    • Multi-Agent Systems for LLMs
    • LLM-as-a-Judge
  • C. Theoretical Analysis (Appendix C)
    • Bi-Directional Mutual Information
    • Internalized Preference Judgement
    • Trajectory of Self-Reflection
    • Objective Decomposition for Unified Multimodal Learning
  • D. Benchmark Details (Appendix D)
    • Data Construction
    • Evaluation Prompt
    • More Results
  • E. Additional Results (Appendix E)
    • Quantitative Results
    • Qualitative Results
    • Failure Cases
@@ Appendix A Experiment Details
  • TIIF: This benchmark evaluates the model’s ability to follow complex prompts across different lengths, specifically categorized into TIIF-S (short) and TIIF-L (long) to measure fine-grained text-to-image alignment. We use the official testmini subset and choose QwenVL2.5-72B bai2025qwen2 as the evaluators.
@@ Appendix B More Related Work
UMMs aim to unify cross-modal understanding and generation, yet a persistent challenge is that strong understanding does not reliably translate into equally strong native generation. Most UMMs follow two main architectural routes. Pure autoregressive models jointly predict text and visual tokens with a unified next-token objective over interleaved sequences, as in Janus-Pro, Emu, and MetaMorph chen2025janus; cui2025emu3; tong2025metamorph. Hybrid designs keep autoregressive modeling for language while relying on diffusion-style synthesis for continuous images, either via integrated modeling within a single backbone (e.g., Show-o and MonoFormer xie2024show; zhao2024monoformer) or through modular routing and sparse experts (e.g., LMFusion, Mixture-of-Transformers, and BAGEL shi2024lmfusion; liang2024mixture; deng2025emerging), with related paradigms such as Diffusion Forcing exploring diffusion-style guidance for interleaved generation chen2024diffusion; huang2025interleaving. Recent work also explores image generation foundation models that systematize text-conditioning and training recipes on diffusion-style backbones. Qwen-Image wu2025qwen adopts a double-stream MMDiT, conditioned on a frozen Qwen2.5-VL encoder and a VAE tokenizer, and uses progressive/curriculum training with multi-task objectives to cover settings such as multilingual text rendering and editing. Z-Image cai2025z proposes a 6B single-stream diffusion Transformer (S3-DiT) and derives a Turbo variant via few-step distillation and reward post-training, focusing on inference efficiency under low sampling steps and text-rendering-related scenarios. Beyond architecture, recent work investigates self-improvement by turning self-produced signals into training objectives yu2025guided; zhou2024calibrated; wang2025unified. For UMMs, SRUM leverages internal understanding as an evaluator to derive fine-grained rewards srum, and UniRL couples generation and understanding via supervised and reinforcement learning unirl. Complementary directions also study data-centric enhancement for vision-language alignment, e.g., ultra-detailed caption generation to enrich training signals for VLMs zeng2025enhancing, and interaction-based learning setups that emphasize perception and reasoning behaviors zeng2025agentic. However, existing self-improvement pipelines often depend on auxiliary components or externally produced dense feedback, as well as task-specific reward shaping, fixed prompt pools, or pre-selected concepts, which can limit scalability and generalization when extending self-improvement to broader UMM settings.
@@
Self-play drives autonomous improvement by pairing self-generated challenges with outcome-driven learning silver2017mastering. In LLMs, this enables zero-data self-evolution: models generate training signals without curated datasets, as exemplified by the uncertainty and self-consistency curricula of R-Zero huang2025r as well as the executor-verified rewards of Absolute Zero zhao2025absolute. Beyond task generation, self-produced evaluation guides preference learning and reasoning, through methods such as self-rewarding feedback yuan2024self, constraint-based optimization zhou2024calibrated, reflective reward learning choi2024self, and process-consistency rewards for long-horizon tasks guo2025r1. In multimodal settings, related efforts incorporate retrieval-augmented reasoning with reinforcement learning to improve understanding over visually rich information sources wang2025vrag. Extending to Unified Multimodal Models (UMMs), their integrated modules naturally enable self-improvement, with the understanding module providing internal multi-scale feedback to guide generation, establishing a promising paradigm for fully model-driven enhancement.
@@
LLM-based multi-agent systems instantiate role-specialized agents to decompose tasks, explore diverse solutions, and cross-check results, supported by orchestration frameworks such as AGENTVERSE chen2024agentverse and debate-style interactions that encourage diverse hypotheses and mutual critique liang2024encouraging. While many systems are primarily deployed at inference time, recent work explores closed-loop training with self-generated signals, including multi-agent training and self-play pipelines motwani2024malt; zhao2025absolute; chen2025multi. Beyond pure language agents, embodied and action-centric extensions seek to improve generalization by structuring the coupling between reasoning and action, as in DualVLA fang2025dualvla. However, empirical analyses show these systems can be brittle and costly, with recurring coordination and verification failures that limit scalability and generalization cemri2025multi. Motivated by these limitations, our work uses lightweight role instantiation within a single unified multimodal model, turns role interactions into self-training signals for improving native multimodal generation, and introduces a cycle-consistency benchmark to test whether gains reflect genuine multimodal understanding rather than task-specific tuning.
@@
Recent work increasingly adopts LLM-as-a-Judge as a scalable alternative to human evaluation for open-ended generation and benchmark construction, where strong LLMs provide pointwise scores or comparative rankings with broader coverage than heuristic metrics at lower annotation cost li2025generation; li2024llms. However, LLM-based judges are not uniformly reliable. Their judgements can be sensitive to prompt phrasing and candidate presentation, and they may exhibit systematic biases and vulnerabilities, including adversarial manipulation raina2024llm; thakur2025judging; li2024llms. These concerns motivate meta-evaluation of judges and evaluation protocols that reduce reliance on fragile or implicit judgement signals feng2025we, as well as targeted benchmarks that stress-test self-critique in tool-calling error scenarios huang2025critictool. In multimodal evaluation, video reasoning benchmarks further broaden coverage beyond static images, including chain-of-thought video reasoning and visual-prompt-based interaction protocols qi2025vcr; zhao2025v2p. In our setting, judge models serve two roles. For T2I generation, we use a VLM-based judge with task-specific rubrics to assess prompt-image alignment and visual fidelity. For T2I2T evaluation, we use the same UMM as an LLM-based judge to verify whether predicted answers match the original instruction and reference answers, enabling structured scoring at scale.

### v2

@@ 1 Introduction
The realization of Artificial General Intelligence (AGI) requires a tight synergy between comprehension and generation, wherein comprehension enables the internalization of knowledge and generation allows its coherent and expressive externalization. By integrating multiple modalities into a shared representational space, Unified Multimodal Models (UMMs) Deng et al. (2025b); Chen et al. (2025b); Xie et al. (2025c) naturally couple comprehension and generation as two complementary phases of a unified cognitive process, supporting both knowledge grounding and coherent reasoning.
@@
Across extensive experiments, we find that our model achieves reliable self-improvement without heuristic reward engineering, curriculum design, or external supervision. Compared with prior self-improvement approaches Jin et al. (2025a) and methods that depend on external guidance, our approach learns from internally generated training signals, generalizes well, and remains stable under out-of-distribution (OOD) conditions. These results support the effectiveness of a fully self-contained learning paradigm.
@@ 2 Related Work
UMMs aim to unify cross-modal understanding and generation, yet strong understanding often fails to yield equally strong native generation. Existing designs fall into two paradigms: pure autoregressive models that jointly predict text and visual tokens over interleaved sequences  Chen et al. (2025e); Cui et al. (2025); Tong et al. (2025)) and hybrid models that combine autoregressive language modeling with diffusion-based image synthesis, either within a unified backbone  Xie et al. (2024); Zhao et al. (2024)) or via modular routing and sparse experts Shi et al. (2024); Liang et al. (2024b); Deng et al. (2025a)), with related guidance schemes such as Diffusion Forcing Chen et al. (2024a). Beyond architecture, self-improvement methods convert self-generated signals into training objectives Yu et al. (2025); Zhou et al. (2024b); Wang et al. (2025b); for UMMs, SRUM derives internal rewards from understanding Jin et al. (2025a), and UniRL jointly optimizes understanding and generation Mao et al. (2025). However, most pipelines depend on auxiliary components or task-specific feedback, limiting scalability and generalization.
@@
Multi-agent systems decompose reasoning through role specialization and interaction, enabling solution diversity and cross-verification, but often incur high coordination cost and brittle verification Chen et al. (2024d); Liang et al. (2024a); Cemri et al. (2025). In parallel, LLM self-improvement converts self-generated tasks and evaluations into training signals, supporting zero-data learning via self-play and self-rewarding mechanisms Silver et al. (2017); Huang et al. (2025a); Zhao et al. (2025a); Yuan et al. (2024). Unified Multimodal Models (UMMs) naturally unify understanding and generation within a single parameter space, making them particularly well-suited for lightweight role instantiation and fully model-driven self-improvement without external supervision.
@@ 3.1 Motivation
Just as a child who associates the word “apple” with the fruit can spontaneously name it upon seeing it, cognitive symmetry Blanco (2018) enables a bidirectional mapping between internal concepts and external expressions. This alignment is reminiscent of escaping Plato’s Cave: true intelligence must move beyond observing surface data to mastering the reciprocal relationship between an appearance and its underlying source.
@@
Figure 3: Results of BAGEL Deng et al. (2025b) and GPT-4o Hurst et al. (2024) on four understanding benchmarks. For Omini-RewardBench Jin et al. (2025b) and MMRB2 Hu et al. (2025), we evaluate the T2I task. Performances are normalized with GPT-4 Achiam et al. (2023) results for better visualization.
@@ 3.3.1 Stage 1: Self Multi-Agent Sampling
LLMs are naturally suited for self-play in multitask settings Radford et al. (2019). For UMMs, interleaved multimodal inputs and functional diversity allow prompting, generation, and judgement to coexist within a shared model, enabling role-conditioned behaviors under different prompts. We leverage this property to functionalize a single UMM into collaborative roles, bridging the comprehension–generation gap through internal synergy.
@@
The proposer is designed to generate a diverse set of challenging prompts for the unified multimodal model, which are subsequently used to produce training images. To this end, inspired by LAION-5B Schuhmann et al. (2022) and COYO-700M Byeon et al. (2022), we partition all T2I task prompts into ten categories and designed fine-grained generation rules for each category. Next, we prompt UMM to generate an initial batch of prompts and act as the judge to select the best candidate for subsequent iterations. Leveraging the strong in-context learning (ICL) capabilities of LLMs Dong et al. (2024), the initial example serves as a few-shot demonstration to guide the generation of subsequent prompts. To further enhance diversity, we introduce a dynamic seeding mechanism. After generating a predefined number of prompts, several examples are sampled from the prompt library for evaluation and then used to construct new demonstrations that guide the next round of prompt generation. Compared with prior approaches that either directly rely on training set Jin et al. (2025a) or employ external models for prompt construction Mao et al. (2025), our method requires no external data and generates more diverse prompts, thereby improving generalization.
@@
The solver is responsible for producing a diverse set of outputs in response to the prompts generated by the proposer. Therefore, we encourage the UMM to generate images under random seeds and different hyperparameters. Following DeepSeek-R1 Guo et al. (2025a), we perform 8 rollouts per prompt to strike a favorable trade-off between sample quality, diversity, and computational efficiency.
@@
Previous work has relied on heuristic reward functions based on keywords Mao et al. (2025) or on powerful external models to provide dense reward maps Jin et al. (2025a). Such reward judges depend heavily on parameter tuning and the performance of external models, which varies across tasks, thereby severely limiting the generalization of self-improvement. As illustrated in Fig. 3, UMMs exhibit strong reward modeling capabilities. Thus, we formulate reward evaluation for all T2I tasks using discrete scores ranging from 0 to 10, following a widely adopted LLM-as-a-judge paradigm Radford et al. (2019); Kim et al. (2023). To further enhance judgement quality, we transfer generation reward models Liu et al. (2025), which have demonstrated strong potential in LLMs, to T2I evaluation. Specifically, we design task-specific rubrics for each category and encourage the model to explicitly articulate its reasoning before producing the final score.
@@ 3.3.2 Stage2: Cognitive Pattern Reconstruction
Through self multi-agent rejection sampling using the Proposer–Solver–Judge pipeline, we obtain a batch of high-quality prompt–image pairs. While these pairs reflect a mapping from abstract conceptual spaces to high-dimensional visual manifolds, directly optimizing this cross-domain alignment remains stochastic and inefficient, often leading to mode collapse Chen et al. (2025a); Zhenyu et al. (2024). To move beyond this "black-box" optimization, we draw inspiration from metacognitive theory Dunlosky and Metcalfe (2008), which identifies monitoring, evaluation, and regulation as the pillars of robust learning. Based on this insight, we propose a tripartite data architecture that reclaims and structures the overlooked trajectories from the self-play cycle. By replaying these latent interactions as explicit caption, judgement, and reflection patterns, we respectively ground abstract concepts in visual features, provide evaluative signals, and encode self-correction processes. This design transforms the previously discarded internal "inner monologue" into a structured supervisory signal, fostering cognitive symmetry without external intervention.
@@
Reflection Inspired by Reflexion Shinn et al. (2023), this pattern introduces iterative regulation to enhance the model’s capacity for self-evolution. Leveraging the Solver’s multiple rollouts $\{I_{1},\dots,I_{n}\}$, we utilize the rewards assigned by the Judge to identify pairs of contrasting quality, specifically selecting a high-reward "winning" image $I^{*}$ and a lower-reward "losing" image $I_{lose}$ from the same prompt. We then construct reflection trajectories formulated as $\pi_{\theta}(I^{*}\mid T,I_{lose},J)$, which explicitly encode the transition from suboptimal states to superior ones. By learning to transform the lower-quality manifestation $I_{lose}$ into its optimized counterpart $I^{*}$, the model internalizes a mechanism for self-correcting generative errors, effectively mitigating mode collapse without the need for external supervision.
@@ 3.4 UniCycle
Based on TIIF Wei et al. (2025), we generate QA pairs to probe instruction-implied attributes grounded in the generated image, extending the original TIIF benchmark from the T2I setting to the Text-to-Image-to-Text (T2I2T) setting. After annotation, we obtain 1,401 TIIF-style instances that cover more than ten task categories and span multiple question formats, including multiple-choice, binary (yes/no), and open-ended questions.
@@ 4.1 Experiment Setup
We adopt BAGEL Deng et al. (2025b) as the base model for our main experiments. The Proposer generates 5,000 prompts, then the Solver rolls out 8 times for each prompt. Training is conducted for 600 steps on 8 NVIDIA H800 GPUs for about 7 hours with a constant learning rate of $1\times 10^{-5}$. Additional details are provided in the Appendix  A.
@@
To validate our method, we compare it against three categories of approaches. First, we consider baseline models, including T2I frameworks: SD3 Medium Esser et al. (2024), FLUX.1-dev Labs (2024) and unified multimodal models: Janus-Pro Chen et al. (2025d), Show-o2 Xie et al. (2025c), BLIP3-o Chen et al. (2025b), UniGen Tian et al. (2025), TwiG Guo et al. (2025c) and T2I-R1 Jiang et al. (2025).
@@
Regarding evaluation, we focus on TIIF Wei et al. (2025), WISE Niu et al. (2025), OneIG-EN Chang et al. (2025), CompBench Kil et al. (2024), DPG Hu et al. (2024), and Geneval Ghosh et al. (2023) to assess generation performance. To evaluate understanding, we further report results on standard benchmarks including MME Fu et al. (2023), MMB Liu et al. (2024b), MMMU-val Yue et al. (2024), MMVP Tong et al. (2024), and MMStar Chen et al. (2024c).
@@ 4.3.2 Model Architecture
We first demonstrate the effectiveness of our method on BAGEL, where understanding and generation components are decoupled. To evaluate its generalization to tightly coupled architectures, we conduct a base model ablation on the purely autoregressive Janus-Pro-7b Chen et al. (2025d). Tab. 4 shows that our method improves Janus by +3.2 on TIIF, +7.0 on WISE, and +4.7 on OneIG-EN, with the most pronounced gain on the knowledge-intensive WISE benchmark. This suggests that the proposed approach enhances knowledge expression by guiding generation through improved understanding, a mechanism that generalizes across different model architectures.
@@ 4.3.3 Scaling Law for UniCorn
Scaling laws guide architectural design and optimization Kaplan et al. (2020); Chen et al. (2024e), but prior methods scale poorly due to reliance on external models or fixed prompts. In contrast, UniCorn achieves scalable self-improvement through unbounded self-sampling and efficient Cognitive Pattern Reconstruction. To explicitly assess scalability, we conduct scale-up experiments by varying the amount of self-generated data across {1k, 5k, 8k, 10k, 20k}.
@@
As shown in Fig. 7, with only 1K training samples, our method already surpasses RecA  (Xie et al., 2025b) on TIIF. As the data scale increases, the model’s generative performance continues to improve, with more pronounced gains on long-prompt generation; notably, with just 5k samples, it outperforms IRG Huang (2025) trained on 30k GPT-4o distilled data as well as the strong closed-source model DALL·E 3 Betker et al. (2023a). These results reveal a favorable scaling regime in which self-generated data alone suffices to drive continual and efficient improvements in generative capability.
@@ 4.4 Analysis
For self-play assessment, we use Qwen3-VL-235B-A22B-Instruct Yang et al. (2025) for data construction (UniCorn*). As shown in Tab. 4, employing stronger proposers/judges yields diminishing returns, where high costs and training time outweigh performance gains. This likely stems from the difficulty of fitting high-entropy teacher distributions, increasing latency without proportional information gain. We then compare UniCorn with four unified models on UniCycle (Judge: Qwen3-235B-A22B Yang et al. (2025)). UniCycle requires both generation and understanding, reducing task bias and evaluating the model’s self-reflection, thus signaling comprehensive multimodal intelligence.
@@ Appendix A Experiment Details
  • TIIF: This benchmark evaluates the model’s ability to follow complex prompts across different lengths, specifically categorized into TIIF-S (short) and TIIF-L (long) to measure fine-grained text-to-image alignment. We use the official testmini subset and choose QwenVL2.5-72B Bai et al. (2025) as the evaluators.
@@ Appendix B More Related Work
UMMs aim to unify cross-modal understanding and generation, yet a persistent challenge is that strong understanding does not reliably translate into equally strong native generation. Most UMMs follow two main architectural routes. Pure autoregressive models jointly predict text and visual tokens with a unified next-token objective over interleaved sequences, as in Janus-Pro, Emu, and MetaMorph Chen et al. (2025e); Cui et al. (2025); Tong et al. (2025). Hybrid designs keep autoregressive modeling for language while relying on diffusion-style synthesis for continuous images, either via integrated modeling within a single backbone (e.g., Show-o and MonoFormer Xie et al. (2024); Zhao et al. (2024)) or through modular routing and sparse experts (e.g., LMFusion, Mixture-of-Transformers, and BAGEL Shi et al. (2024); Liang et al. (2024b); Deng et al. (2025a)), with related paradigms such as Diffusion Forcing exploring diffusion-style guidance for interleaved generation Chen et al. (2024a); Huang et al. (2025c). Recent work also explores image generation foundation models that systematize text-conditioning and training recipes on diffusion-style backbones. Qwen-Image Wu et al. (2025a) adopts a double-stream MMDiT, conditioned on a frozen Qwen2.5-VL encoder and a VAE tokenizer, and uses progressive/curriculum training with multi-task objectives to cover settings such as multilingual text rendering and editing. Z-Image Cai et al. (2025) proposes a 6B single-stream diffusion Transformer (S3-DiT) and derives a Turbo variant via few-step distillation and reward post-training, focusing on inference efficiency under low sampling steps and text-rendering-related scenarios. Beyond architecture, recent work investigates self-improvement by turning self-produced signals into training objectives Yu et al. (2025); Zhou et al. (2024b); Wang et al. (2025b). For UMMs, SRUM leverages internal understanding as an evaluator to derive fine-grained rewards Jin et al. (2025a), and UniRL couples generation and understanding via supervised and reinforcement learning Mao et al. (2025). Complementary directions also study data-centric enhancement for vision-language alignment, e.g., ultra-detailed caption generation to enrich training signals for VLMs Zeng et al. (2025b), and interaction-based learning setups that emphasize perception and reasoning behaviors Zeng et al. (2025a). However, existing self-improvement pipelines often depend on auxiliary components or externally produced dense feedback, as well as task-specific reward shaping, fixed prompt pools, or pre-selected concepts, which can limit scalability and generalization when extending self-improvement to broader UMM settings.
@@
Self-play drives autonomous improvement by pairing self-generated challenges with outcome-driven learning Silver et al. (2017). In LLMs, this enables zero-data self-evolution: models generate training signals without curated datasets, as exemplified by the uncertainty and self-consistency curricula of R-Zero Huang et al. (2025a) as well as the executor-verified rewards of Absolute Zero Zhao et al. (2025a). Beyond task generation, self-produced evaluation guides preference learning and reasoning, through methods such as self-rewarding feedback Yuan et al. (2024), constraint-based optimization Zhou et al. (2024b), reflective reward learning Choi et al. (2024), and process-consistency rewards for long-horizon tasks Guo et al. (2025b). In multimodal settings, related efforts incorporate retrieval-augmented reasoning with reinforcement learning to improve understanding over visually rich information sources Wang et al. (2025a). Extending to Unified Multimodal Models (UMMs), their integrated modules naturally enable self-improvement, with the understanding module providing internal multi-scale feedback to guide generation, establishing a promising paradigm for fully model-driven enhancement.
@@
LLM-based multi-agent systems instantiate role-specialized agents to decompose tasks, explore diverse solutions, and cross-check results, supported by orchestration frameworks such as AGENTVERSE Chen et al. (2024d) and debate-style interactions that encourage diverse hypotheses and mutual critique Liang et al. (2024a). While many systems are primarily deployed at inference time, recent work explores closed-loop training with self-generated signals, including multi-agent training and self-play pipelines Motwani et al. (2024); Zhao et al. (2025a); Chen et al. (2025f). Beyond pure language agents, embodied and action-centric extensions seek to improve generalization by structuring the coupling between reasoning and action, as in DualVLA Fang et al. (2025). However, empirical analyses show these systems can be brittle and costly, with recurring coordination and verification failures that limit scalability and generalization Cemri et al. (2025). Motivated by these limitations, our work uses lightweight role instantiation within a single unified multimodal model, turns role interactions into self-training signals for improving native multimodal generation, and introduces a cycle-consistency benchmark to test whether gains reflect genuine multimodal understanding rather than task-specific tuning.
@@
Recent work increasingly adopts LLM-as-a-Judge as a scalable alternative to human evaluation for open-ended generation and benchmark construction, where strong LLMs provide pointwise scores or comparative rankings with broader coverage than heuristic metrics at lower annotation cost Li et al. (2025, 2024b). However, LLM-based judges are not uniformly reliable. Their judgements can be sensitive to prompt phrasing and candidate presentation, and they may exhibit systematic biases and vulnerabilities, including adversarial manipulation Raina et al. (2024); Thakur et al. (2025); Li et al. (2024b). These concerns motivate meta-evaluation of judges and evaluation protocols that reduce reliance on fragile or implicit judgement signals Feng et al. (2025), as well as targeted benchmarks that stress-test self-critique in tool-calling error scenarios Huang et al. (2025b). In multimodal evaluation, video reasoning benchmarks further broaden coverage beyond static images, including chain-of-thought video reasoning and visual-prompt-based interaction protocols Qi et al. (2025); Zhao et al. (2025c). In our setting, judge models serve two roles. For T2I generation, we use a VLM-based judge with task-specific rubrics to assess prompt-image alignment and visual fidelity. For T2I2T evaluation, we use the same UMM as an LLM-based judge to verify whether predicted answers match the original instruction and reference answers, enabling structured scoring at scale.